OpenStudy (anonymous):

Can someone help set up this problem up? System of equation in 3 variables: The top three mothers day gifts are jewelry, flowers, and gift cards.The total average amounts spent on these gift are $53.42. The Average amount spent on jewelry is $4.40 more than the average amount spent on gift cards. Together, the average amounts spent on flowers and gift cards is $15.58 more than the average amount spent on jewelry. What is the amount spent on each type of gift?

OpenStudy (anonymous):

j + f + g = 53.42 j = g + 4.40 f + g = j + 15.58 Re-write equations 2 and 3 so that the variables are all on the left side.

OpenStudy (anonymous):

Would equations 2 read j - g = 4.4p and equation 3 read f + g -j =15.58?

OpenStudy (anonymous):

yes